# Emilia: A Large-Scale, Extensive, Multilingual, and Diverse Dataset for Speech Generation

### v1

@@ IV-A1 Overview
Using Emilia-Pipe, we construct the Emilia dataset from in-the-wild speech data sourced from a vast collection of video and podcast platforms on the Internet. This data covers various content categories such as talk shows, interviews, debates, sports commentary, and audiobooks, thereby capturing a wide array of real human speaking styles. After processing, the Emilia dataset includes a total of 101,654 hours of multilingual speech data across six different languages.
@@ IV-A2 Dataset Analysis
To validate the quality and diversity of the Emilia dataset, we conduct respective analyses.
@@
To evaluate the quality, we compared Emilia with several existing datasets using DNSMOS scores. This non-intrusive speech quality metric reflects the overall quality of the speech data and is highly correlated with human ratings [37]. Table III presents the speech quality comparison between Emilia and several existing datasets. Emilia achieves a DNSMOS score of 3.26, ranking third among all datasets. The results indicate that, despite being sourced from raw speech data in the wild, after preprocessing, the speech quality of the Emilia dataset is comparable to existing datasets sourced from studio recordings or audiobooks and outperforms the existing datasets sourced from unprocessed in-the-wild speech data.
TABLE III: Quality comparison between Emilia and nine existing datasets. The scores for LJSpeech, AutoPrepWild, AISHELL-3, and LibriTTS are derived from [12]. The score for Libri-Light is computed from its official ”small” subset, and the score for WenetSpeech4TTS is computed from its official “basic” subset. The scores for MLS and Emilia are computed from a randomly sampled 600-hour subset.
Dataset DNSMOS
@@
The Emilia dataset comprises a collection of speech data from a wide range of video and podcast platforms, capturing diverse speaking styles of real human speech. To quantify this diversity, we conducted analyses on both the acoustic and semantic feature spaces, comparing it with the MLS dataset, which is derived from audiobooks and widely used for training speech generation models. Specifically, we randomly selected 5k samples each from the English subset of MLS and Emilia, denoted as Emilia English and MLS English respectively in Fig. 3.
To analyze the diversity of acoustic features, we leveraged a pre-trained WavLM model [38] to extract acoustic representations, capturing various characteristics such as speaker identity, emotion, and prosody. We then applied Principal Component Analysis (PCA) [39] to reduce the dimensionality of these representations to two dimensions. As shown in Fig. 3(a), the Emilia dataset exhibits a broader dispersion compared to MLS, which shows a more compact clustering. The more scattered pattern highlights that the Emilia dataset encompasses a richer coverage of acoustic characteristics compared to the MLS dataset derived from audiobooks.
For the semantic diversity analysis, we employed a pre-trained Sentence-BERT model [40] to generate text representations for the transcripts of each speech sample. Consequently, each speech sample is represented as a 768-dimensional vector based on its textual content, providing a comprehensive approximation of its semantics. Similar to the acoustic analysis, we reduced the dimensionality of the semantic features to two dimensions. As shown in Fig. 3(b), the scatter of textual features indicates that the Emilia dataset covers a wide range of textual content, validating the significant diversity in Emilia’s semantic coverage.
@@ IV-B The Emilia-Large Dataset
Given the significant advancements in large-scale speech generation [23], we are motivated to further scale up the volume of the training dataset for speech generation models to investigate the data scaling laws in speech generation and potentially enhance model performance. In this work, we expand the initial Emilia dataset to an even larger scale, introducing Emilia-Large, which contains a total of 216,313 hours of spontaneous and human-like speech data. Emilia-Large builds upon the same construction methodology as Emilia using our proposed Emilia-Pipe but doubles the total dataset size, making it the largest open-source speech generation dataset to date.
The expansion from Emilia to Emilia-Large primarily uses YODAS2 [41] as the raw data source, termed Emilia-YODAS. The relationship between Emilia and Emilia-Large is illustrated in Fig. 4. YODAS2 is a large-scale (500k hours) real-world speech collection from YouTube videos with CC-BY-3.0 licenses in more than 100 languages. We selectively downloaded and processed data from YODAS2 and other smaller sources in the six languages of the original Emilia dataset.  [^11] Unused data of other languages in YODAS2 can also be seamlessly processed by Emilia-Pipe for speech generation training. Their effectiveness may depend on the ASR model’s performance in each language. We process the raw speech data using our proposed Emilia-Pipe, with only one alteration: changing the DNSMOS filtering threshold for the incremental data to 2.4 to align with [12] and preserve more data. We make the Emilia-YODAS publicly available under the CC-BY-4.0 license.
@@
Emilia-Large (216k hours) Emilia-YODAS ($C$-BY-4.0, 114k hours)Emilia ($C$-BY-NC-4.0, 101k hours)Other small sources
Figure 4: The relationship between Emilia and Emilia-Large: Emilia-Large extends Emilia with additional data, Emilia-YODAS, primarily processed from YODAS2.
@@ V Experiments
In this section, we conduct experiments to address the following evaluation questions (EQs), which are designed to validate the strengths of the Emilia dataset in terms of diversity (EQ1), extensiveness (EQ2), and multilingual utility (EQ3):
  • EQ1: How does the performance of the in-the-wild dataset, Emilia, compare to that of existing audiobook datasets in training speech generation models?
@@
  • EQ3: How effective is the Emilia dataset for training multilingual and crosslingual speech generation models?
V-A Comparison of Audiobook and In-the-Wild Datasets (EQ1)
In this subsection, we conduct experiments to compare audiobook datasets with in-the-wild datasets for text-to-speech (TTS) generation.
@@ V-A Comparison of Audiobook and In-the-Wild Datasets (EQ1)
We implement two SOTA TTS models as baselines: (1) AR+SoundStorm [24]: A two-stage model where a llama-style autoregressive (AR) generative transformer first predicts semantic tokens using text and prompt semantic tokens as input. This is followed by a SoundStorm-based semantic-to-acoustic model [4] that generates acoustic tokens conditioned on the predicted semantic tokens. (2) VoiceBox [3]: A non-autoregressive (NAR) speech generation model that leverages flow matching and in-context learning to predict mel-spectrograms. For comprehensive details on these models, we refer readers to their respective publications.
@@
We evaluate the performance of TTS models trained on two English datasets: the English subset of the Emilia dataset (hereafter referred to as Emilia-En for brevity) and the MLS dataset, a high-quality corpus derived from audiobooks. The Emilia-En dataset consists of approximately 46k hours of English speech, while the MLS dataset contains 45k hours, rendering their sizes comparable.
@@
To ensure a comprehensive evaluation, we employ two distinct evaluation sets: (1) LibriSpeech-Test: This set includes 1,200 speech samples in formal reading styles similar to those in the MLS dataset. (2) Emilia-Test: This set consists of 600 speech samples in spontaneous, human-like speaking styles akin to those in the Emilia dataset. Both sets are unseen by the baseline models during training.
@@
For objective evaluation, we focus on the following aspects. (1) Intelligibility: Measured by the Word Error Rate (WER) of the generated speech’s transcription compared to the target text. For the LibriSpeech-Test, we utilize a fine-tuned HuBERT-Large ASR model. [^12] https://huggingface.co/facebook/hubert-large-ls960-ft For Emilia-Test, we employ the more robust Whisper-Medium [35] for ASR. (2) Similarity: Assessed by Speaker Similarity Score (S-SIM) between the generated speech and the reference speech using the WavLM-TDCNN speaker embedding model. [^13] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification We report the similarity to the original reference speech. (3) Naturalness: Evaluated using the Fréchet Speech Distance (FSD), which measures the distance between the distributions of generated and real samples in a feature space [3]. We compute this metric using emotion2vec features [42] to evaluate the emotional similarity of the speech.
For the subjective evaluation, we randomly select eight samples each from the LibriSpeech-Test and the Emilia-Test evaluation set. Twelve proficient English speakers serve as evaluators. The subjective evaluation includes: (1) Speaker Similarity: We employ the Similarity Mean Opinion Score (SMOS) to assess the speaker similarity of the generated speech to the reference speech. The SMOS scale ranges from 1 to 5, with steps of 0.5. (2) Comparative Naturalness: We use the Comparative Mean Opinion Score (CMOS) to evaluate the comparative naturalness of the generated speech against the reference speech. The CMOS scale ranges from -3 (indicating the generated speech is much worse than the reference speech) to 3 (indicating the generated speech is much better than the reference speech), with steps of 1.
TABLE IV: Objective and subjective evaluation results of TTS models trained with Emilia-En and MLS on LibriSpeech-Test and Emilia-Test evaluation sets. The best results for each model are highlighted in bold.
@@
Table IV summarizes the objective and subjective evaluation results of TTS models trained with the Emilia-En and MLS datasets on the LibriSpeech-Test and Emilia-Test evaluation sets. For the LibriSpeech-Test dataset, the AR+SoundStorm model trained on Emilia-En achieved a lower WER (8.4%) and FSD (24.73) compared to its MLS-trained counterpart, while the VoiceBox model trained on MLS achieved the best WER (6.1%), S-SIM (0.625), and FSD (16.83). On the Emilia-Test dataset, the AR+SoundStorm model trained on Emilia-En outperformed the MLS-trained model across all metrics, including WER (6.6%), S-SIM (0.618), FSD (12.73), CMOS (0.19), and SMOS (3.73). Similarly, the VoiceBox model trained on Emilia-En achieved superior results in WER (7.4%), S-SIM (0.601), and SMOS (3.76) compared to the MLS-trained version.
The result validate that, on the audiobook-style LibriSpeech-Test dataset, models trained on both the Emilia-En and MLS datasets achieve comparable levels of intelligibility, speaker similarity and naturalness. This suggests that the Emilia dataset, despite being derived from raw speech data in the wild, is as effective as high-quality datasets like MLS after processing with our proposed Emilia-Pipe. However, on the Emilia-Test dataset, which includes more spontaneous and human-like speech, training with in-the-wild datasets like Emilia significantly enhances a model’s speech generation performance.
@@
The comparison between the in-the-wild Emilia dataset and audiobook datasets for speech generation tasks reveals that while both types of datasets yield comparable performance in formal, audiobook-style speech, the Emilia dataset significantly outperforms audiobook datasets in generating more spontaneous and human-like speech, showcasing significantly superior performance in cloning diverse speaker timbre and speaking styles.
@@ V-B Data Scaling Law in Speech Generation (EQ2)
Given the extensive nature of the Emilia dataset, we conduct experiments to investigate the impact of dataset size on speech generation performance with a fixed number of model parameters, a.k.a, the data scaling law in speech generation. This section describes our experimental setups and results.
@@
We leverage the TTS baselines, evaluation sets, and objective metrics described in Sec. V-A, progressively increasing the training set size to the following volumes: 5k, 10k, 46k (the total duration of English speech in Emilia-En), 100k, and 134k (the total duration of English speech in Emilia-Large) hours, while recording the corresponding performance changes.
5k10k46k100k134k4444666688885.25.25.25.24.74.74.74.74.44.44.44.44.54.54.54.54.24.24.24.27.27.27.27.25.95.95.95.95.75.75.75.75.75.75.75.75.65.65.65.65.75.75.75.75.25.25.25.24.94.94.94.94.94.94.94.94.94.94.94.97.17.17.17.16.56.56.56.56.36.36.36.36.76.76.76.76.46.46.46.4AR+SoundStorm (L)VoiceBox (L)AR+SoundStorm (E)VoiceBox (E)
Figure 5: WER (%) v.s. train set size. (L) denotes results on LibriSpeech-Test, and (E) denotes results on Emilia-Test.
5k10k46k100k134k0.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.590.590.590.590.60.60.60.60.610.610.610.610.630.630.630.630.620.620.620.620.420.420.420.420.460.460.460.460.60.60.60.60.60.60.60.60.610.610.610.610.590.590.590.590.590.590.590.590.610.610.610.610.640.640.640.640.640.640.640.640.420.420.420.420.440.440.440.440.590.590.590.590.580.580.580.580.610.610.610.61AR+SoundStorm (L)VoiceBox (L)AR+SoundStorm (E)VoiceBox (E)
Figure 6: S-SIM v.s. train set size. (L) denotes results on LibriSpeech-Test, and (E) denotes results on Emilia-Test.
5k10k46k100k134k1414141416161616181818182020202022222222242424242222222221.9721.9721.9721.9720.7120.7120.7120.7120.5820.5820.5820.5820.420.420.420.421.3321.3321.3321.3321.2321.2321.2321.2319.9419.9419.9419.9419.9319.9319.9319.9319.8519.8519.8519.8514.8214.8214.8214.8215.2715.2715.2715.2715.6215.6215.6215.6215.4315.4315.4315.4315.3115.3115.3115.3116.3916.3916.3916.3915.1315.1315.1315.1314.6514.6514.6514.6514.7814.7814.7814.7814.4814.4814.4814.48AR+SoundStorm (L)VoiceBox (L)AR+SoundStorm (E)VoiceBox (E)
Figure 7: FSD v.s. train set size. (L) denotes results on LibriSpeech-Test, and (E) denotes results on Emilia-Test.
Fig. 5 illustrates the WER trends for AR+SoundStorm and VoiceBox across training set sizes. For AR+SoundStorm evaluated on LibriSpeech (L), WER decreases from 5.2% at 5k hours to 4.2% at 134k hours, while VoiceBox (L) shows a reduction from 7.2% to 5.6% over the same range. On the Emilia test set (E), AR+SoundStorm (E) improves from 5.7% to 4.9% WER, and VoiceBox (E) declines from 7.1% to 6.4%. Fig. 6 demonstrates S-SIM scores, with AR+SoundStorm (L) rising from 0.587 to 0.620 and VoiceBox (L) increasing from 0.418 to 0.606 as data scales to 134k hours. For Emilia tests, AR+SoundStorm (E) improves from 0.587 to 0.636, while VoiceBox (E) progresses from 0.422 to 0.612. Fig. 7 reveals FSD trends: AR+SoundStorm (L) decreases from 22.00 to 20.40, VoiceBox (L) from 21.33 to 19.85, AR+SoundStorm (E) from 14.82 to 15.31, and VoiceBox (E) from 16.39 to 14.48.
@@
V-B0b Summary (Answer to EQ2)
Our experimental results reveal the data scaling law in speech generation: performance improves as the dataset size increases, but with diminishing returns. Initially, significant improvements are observed when increasing the dataset size to 46k hours. Beyond 46k hours, performance continues to improve, though become less pronounced and at a slower rate, and tends to converge at around 100k hours. These findings suggest that while larger datasets contribute to better performance, there is a point of diminishing returns where additional data provides only marginal benefits, guiding future research in balancing dataset size and computational resources for optimal speech generation performance. For current TTS models, which typically contain around 0.5–1 B parameters [^14] The size of our baselines also falls within this range., a dataset of approximately 100k hours for each specific language represents the most cost-effective choice.
TABLE V: Experimental results of AR+SoundStorm and VoiceBox for multilingual and crosslingual speech generation. The models were trained on the Emilia-Large dataset. Results for multilingual speech generation are highlighted in grey.
@@
WER 5.9% 5.8% 6.4% 5.9% 6.3% 8.3% 6.5% 7.9% 8.8% 8.6% 8.3% 10.2%
En S-SIM 0.568 0.431 0.452 0.529 0.446 0.443 0.588 0.386 0.458 0.490 0.425 0.442
FSD 24.99 99.40 82.84 26.62 89.40 98.36 24.34 91.29 78.53 68.62 92.54 89.49
WER 5.3% 3.6% 5.2% 5.4% 4.9% 5.7% 8.6% 5.6% 6.7% 5.9% 6.4% 7.0%
Zh S-SIM 0.507 0.511 0.509 0.504 0.516 0.523 0.524 0.557 0.524 0.522 0.543 0.591
FSD 56.15 40.09 56.75 57.10 56.71 52.60 109.67 40.04 58.47 72.47 64.73 57.90
WER 5.3% 5.3% 5.3% 5.2% 5.8% 8.1% 7.0% 6.3% 5.6% 6.9% 7.5% 9.3%
Fr S-SIM 0.596 0.527 0.596 0.596 0.572 0.557 0.565 0.485 0.589 0.582 0.547 0.556
FSD 39.89 66.21 39.88 38.48 51.13 54.41 91.08 80.76 42.38 58.16 63.36 61.51
WER 4.5% 4.5% 4.7% 4.2% 4.8% 6.8% 5.2% 7.4% 6.8% 5.2% 6.9% 8.9%
De S-SIM 0.619 0.545 0.603 0.639 0.596 0.591 0.639 0.519 0.577 0.683 0.538 0.586
FSD 39.96 57.82 44.86 33.16 53.38 55.12 83.37 72.18 54.77 34.41 67.89 67.46
WER 4.6% 4.4% 4.7% 4.5% 4.8% 6.6% 7.4% 5.5% 6.9% 6.7% 6.2% 6.6%
Ja S-SIM 0.622 0.557 0.626 0.618 0.641 0.633 0.556 0.525 0.521 0.557 0.584 0.596
FSD 49.42 68.70 44.67 50.47 44.28 52.19 103.68 76.65 63.55 72.41 44.71 56.34
WER 6.2% 4.1% 6.1% 6.2% 6.2% 6.3% 8.0% 5.6% 7.8% 8.3% 5.6% 6.0%
Ko S-SIM 0.657 0.593 0.665 0.656 0.673 0.673 0.589 0.567 0.545 0.597 0.595 0.648
FSD 36.71 58.85 32.27 37.20 31.95 30.27 86.57 63.49 53.75 57.19 52.85 38.82

### v2

@@ IV-A1 Overview
Using Emilia-Pipe, we construct Emilia from in-the-wild speech data sourced from a vast collection of video and podcast platforms on the Internet. This data covers various speaking styles such as audio-books, drama, interviews, talk shows, and commentary, thereby capturing a wide array of real human speaking styles. After processing, Emilia includes a total of 101,654 hours of multilingual speech data across six different languages.
@@ IV-A2 Dataset Analysis
To validate the quality and diversity of Emilia, we conduct respective analyses.
@@
To evaluate the quality, we compared Emilia with several existing datasets using DNSMOS scores. Table III presents the speech quality comparison between Emilia and several existing datasets. Emilia achieves a DNSMOS score of 3.26, ranking third among all datasets. The results indicate that, despite being sourced from source speech data in-the-wild, after preprocessing, the speech quality of Emilia is comparable to existing datasets sourced from studio recordings or audio-books and outperforms the existing datasets sourced from unprocessed in-the-wild speech data.
TABLE III: DNSMOS Scores of Emilia and ten existing datasets. The scores for LJSpeech, AutoPrepWild, AISHELL-3, LibriTTS, and WenetSpeech, are derived from [11]. The score for Libri-Light is computed from its official “small” subset, and the score for WenetSpeech4TTS is computed from its official “basic” subset. The scores for MLS and Emilia are computed from a randomly sampled 600-hour subset.
Dataset DNSMOS $\uparrow$
@@
TABLE IV: Distribution of speaking style domains in Emilia.
Domain Percentage
Audio-book 9.40%
Commentary 17.75%
Documentary 11.46%
Drama 12.09%
Interview 0.93%
News 2.77%
Reading 15.60%
Talk 13.02%
Variety 6.19%
Others 10.79%
@@
To quantify the diversity of speaking style domains within Emilia, we follow WenetSpeech [20] by classifying the source speech data into ten domains: Audio-book, Commentary, Documentary, Drama, Interview, News, Reading, Talk, Variety, and Others.
This classification is performed by analyzing textual information extracted from source URLs and metadata (such as hashtags) using Gemma-3-12B-it. [^11] https://huggingface.co/google/gemma-3-12b-it As shown in Table IV, Emilia demonstrates a well-balanced distribution of speaking styles across most domains
To illustrate the diversity advantage of Emilia over traditional audio-book datasets, we perform a comparative analysis between Emilia and the MLS dataset in both acoustic and semantic feature spaces. Specifically, we randomly select 5,000 samples each from the English subsets of Emilia and MLS, denoted as Emilia English and MLS English.
For acoustic features, we leverage a pre-trained WavLM model [35] to extract 768-dimensional embeddings for each speech sample, which captures various acoustic attributes such as speaker identity, emotion, and prosody. These high-dimensional features are then reduced to two dimensions using Principal Component Analysis (PCA) [36]. As shown in Fig. 3(a), Emilia displays a wider distribution compared to the more compact clustering of MLS, highlighting Emilia’s richer coverage in acoustic characteristics.
For semantic diversity, we employ a pre-trained Sentence-BERT model [37] to generate 768-dimensional text embeddings from the transcripts of each sample. These embeddings capture the semantic content of these samples. Similar to above, we reduce the dimensions of these embeddings to two using PCA. Fig. 3(b) demonstrates that Emilia spans a broader semantic coverage than MLS.
@@ IV-B The Emilia-Large Dataset
Given the significant advancements in large-scale speech generation [23], we are motivated to further scale up the volume of the training dataset for speech generation models to investigate the data scaling laws in speech generation and potentially enhance model performance. In this work, we expand the initial Emilia dataset to an even larger scale, introducing Emilia-Large, which contains a total of 216,313 hours of spontaneous and human-like speech data. Emilia-Large builds upon the same construction methodology as Emilia using our proposed Emilia-Pipe but doubles the total dataset size.
The expansion from Emilia to Emilia-Large primarily uses YODAS2 [38] as the data source, termed Emilia-YODAS. The relationship between Emilia and Emilia-Large is illustrated in Fig. 4. YODAS2 is a large-scale (500k hours) real-world speech collection from YouTube videos with CC-BY-3.0 licenses in more than 100 languages. We selectively downloaded and processed data from YODAS2 and other smaller sources in the six languages of the original Emilia dataset. [^12] Unused data of other languages in YODAS2 can also be seamlessly processed by Emilia-Pipe for speech generation training. Their effectiveness may depend on the ASR model’s performance in each language. We process the source speech data using our proposed Emilia-Pipe, with only one alteration: changing the DNSMOS filtering threshold for the incremental data to 2.4 to align with [11] and preserve more data. We make the Emilia-YODAS publicly available under the CC-BY-4.0 license.
@@
Emilia-Large (216k hours) Emilia-YODAS ($C$-BY-4.0, 114k hours)Emilia ($C$-BY-NC-4.0, 101k hours)Other sources
Figure 4: The relationship between Emilia and Emilia-Large. Emilia-Large is an extended version of Emilia, incorporating additional data primarily processed from YODAS2.
@@ V Experiments
In this section, we conduct experiments to address the following evaluation questions (EQs) to validate the strengths of Emilia in terms of diversity (EQ1), extensiveness (EQ2), and multilingual utility (EQ3):
  • EQ1: Is Emilia more effective than existing audio-book datasets for training spontaneous and human-like speech generation model?
@@
  • EQ3: How effective is Emilia for training multilingual and crosslingual speech generation models?
V-A Comparison of Audio-book and In-the-Wild Datasets (EQ1)
To address EQ1, we compare the performance of TTS models trained on existing audio-book datasets to the performance of models trained on Emilia.
@@ V-A Comparison of Audio-book and In-the-Wild Datasets (EQ1)
We implement two SOTA TTS models as baselines: (1) AR+SoundStorm [39]: A two-stage model where a LLaMA-style autoregressive (AR) generative transformer first predicts semantic tokens using text and prompt semantic tokens as input. Then a SoundStorm-based semantic-to-acoustic model [4] that generates acoustic tokens conditioned on the predicted semantic tokens. (2) VoiceBox [3]: A non-autoregressive (NAR) speech generation model that leverages flow matching to predict mel-spectrograms. For comprehensive details on these models, we refer readers to their respective publications.
@@
We evaluate the performance of TTS models trained on two English datasets: the English subset of Emilia (hereafter referred to as Emilia-En for brevity) and the MLS dataset, a high-quality corpus derived from audio-books. The Emilia-En dataset consists of approximately 46k hours of English speech, while the MLS dataset contains 45k hours, rendering their sizes comparable.
@@
To ensure a comprehensive evaluation, we employ two evaluation sets in two domains: (1) LibriSpeech-Test: This evaluation set includes 1,200 speech samples in formal reading styles similar to those in the MLS dataset. (2) Emilia-Test: This evaluation set consists of 600 speech samples in spontaneous, human-like speaking styles akin to those in Emilia. Both evaluation sets are unseen by the baseline models during training.
@@
For objective evaluation, we focus on the following aspects. (1) Intelligibility: This is measured by the Word Error Rate (WER) between the generated speech transcription and the target text. For LibriSpeech-Test, we use a fine-tuned HuBERT-Large ASR model [^13] https://huggingface.co/facebook/hubert-large-ls960-ft to transcribe the generated speech. For Emilia-Test, we use the Whisper-Medium model [31], which offers greater robustness for in-the-wild speech. (2) Similarity: This is measured by Speaker Similarity Score (S-SIM) between the generated speech and the reference speech using the WavLM-TDCNN speaker embedding model. [^14] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification (3) Naturalness: This is measured using the Fréchet Speech Distance (FSD), which quantifies the distance between the distributions of generated and real speech samples in a feature space [3]. We employ the SOTA emotion representation model, emotion2vec [40], to compute FSD and evaluate the emotional naturalness of the generated speech.
For the subjective evaluation, we randomly select sixteen samples, eight from the LibriSpeech-Test and eight from the Emilia-Test evaluation set. Twelve proficient English speakers serve as evaluators. The subjective evaluation includes: (1) Speaker Similarity: We employ the Similarity Mean Opinion Score (SMOS) to assess the speaker similarity of the generated speech to the reference speech. The SMOS scale ranges from 1 to 5, with steps of 0.5. (2) Comparative Naturalness: We use the Comparative Mean Opinion Score (CMOS) to evaluate the comparative naturalness of the generated speech against the reference speech. The CMOS scale ranges from -3 (indicating the generated speech is much worse than the reference speech) to 3 (indicating the generated speech is much better than the reference speech), with steps of 1.
TABLE V: Objective and subjective evaluation results of TTS models trained on Emilia-En and MLS on LibriSpeech-Test and Emilia-Test evaluation sets. The best results for each model are highlighted in bold.
@@
Table V summarizes the objective and subjective evaluation results of TTS models trained on the Emilia-En and MLS datasets on the LibriSpeech-Test and Emilia-Test evaluation sets.
On LibriSpeech-Test, the AR+SoundStorm model trained on Emilia-En achieved a lower WER (8.4%) and FSD (24.73) compared to its MLS-trained counterpart, while the VoiceBox model trained on MLS achieved the best WER (6.1%), S-SIM (0.625), and FSD (16.83). On the Emilia-Test, the AR+SoundStorm model trained on Emilia-En outperformed the MLS-trained model across all metrics, including WER (6.6%), S-SIM (0.618), FSD (12.73), CMOS (0.19), and SMOS (3.73). Similarly, the Emilia-En trained VoiceBox achieved superior results in WER (7.4%), S-SIM (0.601), and SMOS (3.76) compared to the MLS-trained version.
The results validate that, on the formal reading style LibriSpeech-Test evaluation set, models trained on both the Emilia-En and MLS datasets achieve comparable levels of intelligibility, speaker similarity, and naturalness. This suggests that Emilia, despite being derived from source speech data in-the-wild, is as effective as high-quality datasets like MLS after processing with our proposed Emilia-Pipe. However, on the Emilia-Test evaluation set, which includes more spontaneous and human-like speech, training on in-the-wild datasets like Emilia significantly enhances the performance of speech generation models.
@@
The comparison between the in-the-wild Emilia dataset and audio-book MLS dataset for speech generation tasks reveals that while both types of datasets yield comparable performance in formal, audio-book-style speech, Emilia significantly outperforms audio-book datasets in generating more spontaneous and human-like speech, showcasing significantly superior performance in cloning diverse speaker timbre and speaking styles.
@@ V-B Data Scaling Law in Speech Generation (EQ2)
To address EQ2, we conduct experiments to investigate the impact of dataset size on speech generation performance with a fixed number of model parameters, i.e., the data scaling law in speech generation.
@@
We leverage the baseline models, evaluation sets, and objective metrics described in Sec. V-A. To assess performance scaling, we progressively increase the training set size to the following amounts: 5k, 10k, 46k (the total duration of English speech in Emilia-En), 100k, and 134k hours (the total duration in Emilia-Large). At each stage, we record the corresponding changes in model performance.
V-B0b Results and Discussions
Fig. 5(a) shows the WER trends for AR+SoundStorm and VoiceBox as training set size increases. On LibriSpeech-Test (L), both models exhibit a decrease in WER; for instance, WER of AR+SoundStorm drops from 5.2% at 5k hours to 4.2% at 134k hours. Similarly, on the Emilia test set (E), the WER of AR+SoundStorm decreases from 5.7% to 4.9%, while the WER of VoiceBox declines from 7.1% to 6.4%.
Fig. 5(b) illustrates the S-SIM trends. On LibriSpeech-Test, AR+SoundStorm improves from 0.587 to 0.620, while VoiceBox rises from 0.418 to 0.606 as the dataset scales to 134k hours. On Emilia-Test, AR+SoundStorm increases from 0.587 to 0.636, and VoiceBox advances from 0.422 to 0.612.
Fig. 5(c) shows the FSD trends. On LibriSpeech-Test, AR+SoundStorm decreases from 22.00 to 20.40, and VoiceBox decreases from 21.33 to 19.85. On Emilia-Test, AR+SoundStorm changes from 14.82 to 15.31, while VoiceBox decreases from 16.39 to 14.48.
@@
5k10k46k100k134k4466885.25.24.74.74.44.44.54.54.24.27.27.25.95.95.75.75.75.75.65.65.75.75.25.24.94.94.94.94.94.97.17.16.56.56.36.36.76.76.46.4AR+SoundStorm (L)VoiceBox (L)AR+SoundStorm (E)VoiceBox (E)
(a) WER (%)
5k10k46k100k134k0.40.40.50.50.60.60.70.70.80.80.590.590.60.60.610.610.630.630.620.620.420.420.460.460.60.60.60.60.610.610.590.590.590.590.610.610.640.640.640.640.420.420.440.440.590.590.580.580.610.61AR+SoundStorm (L)VoiceBox (L)AR+SoundStorm (E)VoiceBox (E)
(b) S-SIM
5k10k46k100k134k141416161818202022222424222221.9721.9720.7120.7120.5820.5820.420.421.3321.3321.2321.2319.9419.9419.9319.9319.8519.8514.8214.8215.2715.2715.6215.6215.4315.4315.3115.3116.3916.3915.1315.1314.6514.6514.7814.7814.4814.48AR+SoundStorm (L)VoiceBox (L)AR+SoundStorm (E)VoiceBox (E)
(c) FSD
Figure 5: Model performance vs. training set size on LibriSpeech-Test (L) and Emilia-Test (E).
V-B0c Summary (Answer to EQ2)
Our experimental results demonstrate a data scaling law in speech generation: as dataset size increases, performance improves, but with diminishing returns. Significant gains are observed when scaling up to 46k hours; beyond this point, improvements continue but become less substantial and tend to plateau around 100k hours. This insight can help guide future research in balancing dataset size and computational resources for optimal speech generation. For TTS models containing around 0.5–1 billion parameters, [^15] The size of our baselines also falls within this range. a dataset of approximately 100k hours per language appears to be the most cost-effective choice.
TABLE VI: Experimental results of AR+SoundStorm and VoiceBox for multilingual and crosslingual speech generation. The models were trained on the Emilia-Large dataset. Results for multilingual speech generation are highlighted in gray.
@@
WER $\downarrow$ 5.9% 5.8% 6.4% 5.9% 6.3% 8.3% 6.5% 7.9% 8.8% 8.6% 8.3% 10.2%
En S-SIM $\uparrow$ 0.568 0.431 0.452 0.529 0.446 0.443 0.588 0.386 0.458 0.490 0.425 0.442
FSD $\downarrow$ 24.99 99.40 82.84 26.62 89.40 98.36 24.34 91.29 78.53 68.62 92.54 89.49
WER $\downarrow$ 5.3% 3.6% 5.2% 5.4% 4.9% 5.7% 8.6% 5.6% 6.7% 5.9% 6.4% 7.0%
Zh S-SIM $\uparrow$ 0.507 0.511 0.509 0.504 0.516 0.523 0.524 0.557 0.524 0.522 0.543 0.591
FSD $\downarrow$ 56.15 40.09 56.75 57.10 56.71 52.60 109.67 40.04 58.47 72.47 64.73 57.90
WER $\downarrow$ 5.3% 5.3% 5.3% 5.2% 5.8% 8.1% 7.0% 6.3% 5.6% 6.9% 7.5% 9.3%
Fr S-SIM $\uparrow$ 0.596 0.527 0.596 0.596 0.572 0.557 0.565 0.485 0.589 0.582 0.547 0.556
FSD $\downarrow$ 39.89 66.21 39.88 38.48 51.13 54.41 91.08 80.76 42.38 58.16 63.36 61.51
WER $\downarrow$ 4.5% 4.5% 4.7% 4.2% 4.8% 6.8% 5.2% 7.4% 6.8% 5.2% 6.9% 8.9%
De S-SIM $\uparrow$ 0.619 0.545 0.603 0.639 0.596 0.591 0.639 0.519 0.577 0.683 0.538 0.586
FSD $\downarrow$ 39.96 57.82 44.86 33.16 53.38 55.12 83.37 72.18 54.77 34.41 67.89 67.46
WER $\downarrow$ 4.6% 4.4% 4.7% 4.5% 4.8% 6.6% 7.4% 5.5% 6.9% 6.7% 6.2% 6.6%
Ja S-SIM $\uparrow$ 0.622 0.557 0.626 0.618 0.641 0.633 0.556 0.525 0.521 0.557 0.584 0.596
FSD $\downarrow$ 49.42 68.70 44.67 50.47 44.28 52.19 103.68 76.65 63.55 72.41 44.71 56.34
WER $\downarrow$ 6.2% 4.1% 6.1% 6.2% 6.2% 6.3% 8.0% 5.6% 7.8% 8.3% 5.6% 6.0%
Ko S-SIM $\uparrow$ 0.657 0.593 0.665 0.656 0.673 0.673 0.589 0.567 0.545 0.597 0.595 0.648
FSD $\downarrow$ 36.71 58.85 32.27 37.20 31.95 30.27 86.57 63.49 53.75 57.19 52.85 38.82